# Safe-Control: A Safety Patch for Mitigating Unsafe Content in Text-to-Image Generation Models

### v1

@@ Abstract
Index Terms:
Text-to-Image, safety patch, defense mechanisms.
I Introduction
@@
II Background
II-A Unsafe Content Generation in T2I Models
@@
II-B Safety Mechanisms for T2I Models
@@ II-C2 Defender
III Design of Safe-Control
III-A Overview
@@
The figure left side shows the Training Data Generation process, where we construct a multi-modal dataset by pairing unsafe prompts (e.g., “a naked female”) with corresponding safe images (see Section III-B). Safe prompts are generated from unsafe prompts using a Large Language Model (LLM) and a crafted prompt template, guided by the OpenAI content policy [22] and the definition of unsafe content from Gebru et al. [23]. These safe prompts are fed into a third-party text-to-image model to produce candidate images, which are then filtered using semantic tools and content classifiers to ensure safety. Training Safe-Control on this dataset allows it to learn conditional safety signals, enabling the transformation of unsafe images generated from unsafe prompts into safe ones.
The figure right side shows the Safe-Control training framework. Safe-Control adopts its network architecture from ControlNet [24], using a trainable copy of the target T2I model’s encoding layers as the backbone of the safety patch. Additionally, Safe-Control employs zero convolution connections between the trainable copy and the target T2I model. To enhance performance and safety, Safe-Control implements two key strategies: First, it locks the parameters of the target T2I model, training only the external safety patch. This preserves the model’s ability to generate high-quality images. Second, Safe-Control is trained on both a multi-modal dataset and a benign dataset containing safe prompts and images. This ensures that only conditional safety signals are injected, protecting the model from harmful noise without compromising its utility. Furthermore, Since visual conditions cannot effectively represent “safety,” directly applying ControlNet for safety enforcement is challenging. To address this, we replace ControlNet’s visual condition with safety modification instructions for input prompts, serving as the safety condition to guide the T2I model’s generation process. Additionally, we introduce a safety condition mapping module to handle non-visual safety conditions. See details in Section III-C.
TABLE I: Example pairs of unsafe prompts and their corresponding LLM-generated safe prompts.
@@
III-B Training Data Generation
@@
Generating Safe Prompts. To ensure reliable safe prompts, we use LLMs to generate multiple candidate prompts for each unsafe prompt rather than a single substitution. We define unsafe scenarios based on OpenAI’s content policy [22] and Gebru et al. [23], rather than relying solely on the LLM’s inherent understanding. This approach provides clear context guidance and standardizes unsafe content criteria across the dataset. Fig. 3 presents the prompt template, and Table I lists example prompt pairs.
@@
III-C Finding Safety Condition
ControlNet is designed to regulate content generation based on explicit visual conditions such as posture or depth. However, “safety” is difficult to define as a unified visual condition. Instead, we leverage textual descriptions to enforce safety. For example, for an unsafe prompt like “a naked woman,” the modification instruction could be “Put clothes on the person in the image.” For a neutral prompt like “a monkey,” no modification is needed. Similarly, concepts related to hate or violence can be managed through specific instructions. We introduce safety modification instructions as a control mechanism for T2I models, replacing traditional visual conditions. To achieve this, we use an LLM to assess the prompt’s safety and generate corresponding modification instructions. The prompt templates follow those in Section III-B.
III-D Training Safe-Control
@@ IV Experiment
IV-A Experimental Settings
IV-A1 Datasets
@@ IV-A2 Baselines
IV-A3 Text-To-Image Generation Models
@@ IV-A4 Metrics
IV-B RQ1: Effectiveness in Mitigating Unsafe Generation
@@
TABLE II: The nudity probability of images generated by various T2I models, both with and without our proposed $\textit{Safe-Control}_{nudity}$ patch. The values marked in blue indicate the ratio of nudity content erased by $\textit{Safe-Control}_{nudity}$ relative to the original T2I models. A decrement is represented by “-,” whereas “+” signifies an increment.
@@
TABLE III: The unsafe probability of images generated by various T2I models, both with and without $\textit{Safe-Control}_{multiple}$ patch. The values marked in blue indicate the ratio of unsafe content erased by $\textit{Safe-Control}_{multiple}$ relative to the original T2I models. A decrement is represented by “-,” whereas “+” signifies an increment.
@@
To ensure the comprehensiveness of the evaluation, we selected three different types of prompt datasets. Table II shows the nudity probability of images generated by various T2I models, both with and without our proposed $\textit{Safe-Control}_{nudity}$ patch. Note that $\textit{Safe-Control}_{nudity}$ is a specifically trained safety patch designed to mitigate the generation of nude content. The blue-marked values in the table indicate the ratio of nudity content erased by $\textit{Safe-Control}_{nudity}$ relative to the original T2I models.
Table III delineates the unsafe probability of images generated by various T2I models, both with and without the unified safety patch $\textit{Safe-Control}_{multiple}$. The values marked in blue indicate the ratio of unsafe content erased by $\textit{Safe-Control}_{multiple}$ relative to the original T2I models. We have three observations. First, similarly to the observations made on the nudity content reduction, one can observe in Table III that the probability of all types of unsafe contents markedly decreases when enhanced with our proposed safety patch. For example, for any type of unsafe content, the average unsafe probability reduction achieved by $\textit{Safe-Control}_{multiple}$ across seven T2I models is close to or exceeds 50%, with the highest even reaching up to 70%. Second, $\textit{Safe-Control}_{multiple}$ demonstrates potent defensive capability across seven types of T2I models, managing to remove an average of 61% of unsafe content, thus showing prominent transferability. Third, $\textit{Safe-Control}_{multiple}$ excels at mitigating the generation of sexual content, being able to reduce the generation probability of such unsafe content to no more than 7% in five models. This is primarily because the harm caused by sexual content is relatively high. Therefore, in setting up safety patch combinations, we adopted a strategy of granting $\textit{Safe-Control}_{multiple}$ a higher weight to suppress the generation of such unsafe content. For practical applications, users can adjust the weight of each patch according to their needs, thereby designing their own safety patch.
TABLE IV: The performance of $\textit{Safe-Control}_{nudity}$ in maintaining benign generation for prompts from COCO 2017 val, and its comparison with the original T2I models.
@@
Benign Generation Preservation. To evaluate the ability of preserving benign generation, we discuss the overall impact of Safe-Control on image fidelity and text alignment of original T2I models. Ideally, the approach should have no adverse effect on either, especially on benign images. In line with previous research [10], we employ a benign prompt dataset, MS COCO 2017, and report the FID, LPIPS, and CLIP scores of images generated by various T2I models, both with and without $\textit{Safe-Control}_{nudity}$ patch in Table IV. We obtain three key observations from Table IV.
@@
TABLE V: The nudity removal performance of our proposed $\textit{Safe-Control}_{nudity}$ and baselines compared to original Text-to-Image generative models across different adversarial prompt datasets.
@@ IV-C RQ2: Performance Comparison with Baselines
Table V shows the performance of our proposed defense and baselines on nudity content reduction across three different prompt datasets. With respect to the “<country > body” prompt dataset, we produced an aggregate of 2000 images using the prompt “<country > body” associated with the top-50 GDP countries (20 images for each country). Based on the results across three different prompt datasets, $\textit{Safe-Control}_{nudity}$ has a more significant effect in mitigating nudity content generation than baselines, with the highest average nudity removal rate of 94%. In baselines, only the ESD method is comparable to ours in effectiveness, while the average nudity removal rate of other baselines is over 6% lower than ours.
In addition, we found that the effectiveness of the SLD method was not ideal, primarily because we chose the SLD-medium configuration. When using the stronger SLD-strong and SLD-max configurations, the average nudity removal rates reached 73% and 85%, respectively. However, these two configurations significantly degraded the quality and text alignment of benign generation, as shown in Table VII.
Table VI delineates the performance of our proposed $\textit{Safe-Control}_{multiple}$ and baselines on multiple unsafe content reduction on I2P prompt dataset. First, overall, $\textit{Safe-Control}_{multiple}$ almost reduced the probability of generating unsafe content from all types of unsafe prompts to below 10%, representing an overall improvement of 4% compared to the baselines. Second, due to the potential impact of distributional shifts on the filters used during in data censorship or inference-stage filtering, external defenses (such as SD v2.1 and safety filter) can only reduce the probability of generating unsafe content in T2I models by less than 20%, indicating that their generalization capabilities are insufficient.
TABLE VI: The performance of our proposed $\textit{Safe-Control}_{multiple}$ and baselines on multiple unsafe content reduction on I2P prompt dataset.
@@
Table VII displays the performance of $\textit{Safe-Control}_{multiple}$ in maintaining benign generation for prompts from MS COCO 2017 val, and its comparison with baselines. Note that the downward arrow $\downarrow$ indicates that a lower score is more desirable, while the upward arrow $\uparrow$ suggests that a higher score is preferable. We obtain two key observations from Table IV.
TABLE VII: The performance of $\textit{Safe-Control}_{multiple}$ in maintaining benign generation for prompts from COCO 2017 val, and its comparison with baselines. Note that the downward arrow $\downarrow$ indicates that a lower score is more desirable, while the upward arrow $\uparrow$ suggests that a higher score is preferable.
@@
IV-D RQ3: Robustness Against Malicious Attacks
@@
Results. Table VIII shows the robustness evaluation of our proposed $\textit{Safe-Control}_{multiple}$ method compared to several baseline approaches—Original SD, SLD, InterpreteDiffusion, ESD, and SAFEGEN—under two malicious attacks. The evaluation metrics assess the probability of unsafe content in images generated by T2I models equipped with different defenses.
@@
TABLE VIII: The robustness of our proposed $\textit{Safe-Control}_{multiple}$ and baselines against latest malicious attacks. We report the unsafe probability of images being generated by two attacks.

### v2

@@ Abstract
Introduction
@@
Background
Unsafe Content Generation in T2I Models
@@
Safety Mechanisms for T2I Models
@@
Design of Safe-Control
Overview
@@
The figure left side shows the Training Data Generation process, where we construct a multi-modal dataset by pairing unsafe prompts (e.g., “a naked female”) with corresponding safe images (see Section 3.2). Safe prompts are generated from unsafe prompts using a Large Language Model (LLM) and a crafted prompt template, guided by the OpenAI content policy [5] and the definition of unsafe content from Gebru et al. [10]. These safe prompts are fed into a third-party text-to-image model to produce candidate images, which are then filtered using semantic tools and content classifiers to ensure safety. Training Safe-Control on this dataset allows it to learn conditional safety signals, enabling the transformation of unsafe images generated from unsafe prompts into safe ones.
The figure right side shows the Safe-Control training framework. Safe-Control adopts its network architecture from ControlNet [40], using a trainable copy of the target T2I model’s encoding layers as the backbone of the safety patch. Additionally, Safe-Control employs zero convolution connections between the trainable copy and the target T2I model. To enhance performance and safety, Safe-Control implements two key strategies: First, it locks the parameters of the target T2I model, training only the external safety patch. This preserves the model’s ability to generate high-quality images. Second, Safe-Control is trained on both a multi-modal dataset and a benign dataset containing safe prompts and images. This ensures that only conditional safety signals are injected, protecting the model from harmful noise without compromising its utility. Furthermore, Since visual conditions cannot effectively represent “safety,” directly applying ControlNet for safety enforcement is challenging. To address this, we replace ControlNet’s visual condition with safety modification instructions for input prompts, serving as the safety condition to guide the T2I model’s generation process. Additionally, we introduce a safety condition mapping module to handle non-visual safety conditions. See details in Section 3.3.
Table 1: Example pairs of unsafe prompts and their corresponding LLM-generated safe prompts.
@@
Training Data Generation
@@
Generating Safe Prompts. To ensure reliable safe prompts, we use LLMs to generate multiple candidate prompts for each unsafe prompt rather than a single substitution. We define unsafe scenarios based on OpenAI’s content policy [5] and Gebru et al. [10], rather than relying solely on the LLM’s inherent understanding. This approach provides clear context guidance and standardizes unsafe content criteria across the dataset. Fig. 3 presents the prompt template, and Table 1 lists example prompt pairs.
@@
Finding Safety Condition
ControlNet is designed to regulate content generation based on explicit visual conditions such as posture or depth. However, “safety” is difficult to define as a unified visual condition. Instead, we leverage textual descriptions to enforce safety. For example, for an unsafe prompt like “a naked woman,” the modification instruction could be “Put clothes on the person in the image.” For a neutral prompt like “a monkey,” no modification is needed. Similarly, concepts related to hate or violence can be managed through specific instructions. We introduce safety modification instructions as a control mechanism for T2I models, replacing traditional visual conditions. To achieve this, we use an LLM to assess the prompt’s safety and generate corresponding modification instructions. The prompt templates follow those in Section 3.2.
Training Safe-Control
@@
Experimental Settings
Datasets
@@
Text-To-Image Generation Models
@@
RQ1: Effectiveness in Mitigating Unsafe Generation
@@
Table 2: The nudity probability of images generated by various T2I models, both with and without our proposed $\textit{Safe-Control}_{nudity}$ patch. The values marked in blue indicate the ratio of nudity content erased by $\textit{Safe-Control}_{nudity}$ relative to the original T2I models. A decrement is represented by “-,” whereas “+” signifies an increment.
@@
Table 3: The unsafe probability of images generated by various T2I models, both with and without $\textit{Safe-Control}_{multiple}$ patch. The values marked in blue indicate the ratio of unsafe content erased by $\textit{Safe-Control}_{multiple}$ relative to the original T2I models. A decrement is represented by “-,” whereas “+” signifies an increment.
@@
Table 4: The performance of $\textit{Safe-Control}_{nudity}$ in maintaining benign generation for prompts from COCO 2017 val, and its comparison with the original T2I models.
@@
To ensure the comprehensiveness of the evaluation, we selected three different types of prompt datasets. Table 2 shows the nudity probability of images generated by various T2I models, both with and without our proposed $\textit{Safe-Control}_{nudity}$ patch. Note that $\textit{Safe-Control}_{nudity}$ is a specifically trained safety patch designed to mitigate the generation of nude content. The blue-marked values in the table indicate the ratio of nudity content erased by $\textit{Safe-Control}_{nudity}$ relative to the original T2I models.
Table 3 delineates the unsafe probability of images generated by various T2I models, both with and without the unified safety patch $\textit{Safe-Control}_{multiple}$. The values marked in blue indicate the ratio of unsafe content erased by $\textit{Safe-Control}_{multiple}$ relative to the original T2I models. We have three observations. First, similarly to the observations made on the nudity content reduction, one can observe in Table 3 that the probability of all types of unsafe contents markedly decreases when enhanced with our proposed safety patch. For example, for any type of unsafe content, the average unsafe probability reduction achieved by $\textit{Safe-Control}_{multiple}$ across seven T2I models is close to or exceeds 50%, with the highest even reaching up to 70%. Second, $\textit{Safe-Control}_{multiple}$ demonstrates potent defensive capability across seven types of T2I models, managing to remove an average of 61% of unsafe content, thus showing prominent transferability. Third, $\textit{Safe-Control}_{multiple}$ excels at mitigating the generation of sexual content, being able to reduce the generation probability of such unsafe content to no more than 7% in five models. This is primarily because the harm caused by sexual content is relatively high. Therefore, in setting up safety patch combinations, we adopted a strategy of granting $\textit{Safe-Control}_{multiple}$ a higher weight to suppress the generation of such unsafe content. For practical applications, users can adjust the weight of each patch according to their needs, thereby designing their own safety patch.
Benign Generation Preservation. To evaluate the ability of preserving benign generation, we discuss the overall impact of Safe-Control on image fidelity and text alignment of original T2I models. Ideally, the approach should have no adverse effect on either, especially on benign images. In line with previous research [18], we employ a benign prompt dataset, MS COCO 2017, and report the FID, LPIPS, and CLIP scores of images generated by various T2I models, both with and without $\textit{Safe-Control}_{nudity}$ patch in Table 4. We obtain three key observations from Table 4.
@@
RQ2: Performance Comparison with Baselines
Table 5: The nudity removal performance of our proposed $\textit{Safe-Control}_{nudity}$ and baselines compared to original Text-to-Image generative models across different adversarial prompt datasets.
@@
Table 5 shows the performance of our proposed defense and baselines on nudity content reduction across three different prompt datasets. With respect to the “<country > body" prompt dataset, we produced an aggregate of 2000 images using the prompt “<country > body" associated with the top-50 GDP countries (20 images for each country). Based on the results across three different prompt datasets, $\textit{Safe-Control}_{nudity}$ has a more significant effect in mitigating nudity content generation than baselines, with the highest average nudity removal rate of 94%. In baselines, only the ESD method is comparable to ours in effectiveness, while the average nudity removal rate of other baselines is over 6% lower than ours.
In addition, we found that the effectiveness of the SLD method was not ideal, primarily because we chose the SLD-medium configuration. When using the stronger SLD-strong and SLD-max configurations, the average nudity removal rates reached 73% and 85%, respectively. However, these two configurations significantly degraded the quality and text alignment of benign generation, as shown in Table 7.
Table 6: The performance of our proposed $\textit{Safe-Control}_{multiple}$ and baselines on multiple unsafe content reduction on I2P prompt dataset.
@@
Table 6 delineates the performance of our proposed $\textit{Safe-Control}_{multiple}$ and baselines on multiple unsafe content reduction on I2P prompt dataset. First, overall, $\textit{Safe-Control}_{multiple}$ almost reduced the probability of generating unsafe content from all types of unsafe prompts to below 10%, representing an overall improvement of 4% compared to the baselines. Second, due to the potential impact of distributional shifts on the filters used during in data censorship or inference-stage filtering, external defenses (such as SD v2.1 and safety filter) can only reduce the probability of generating unsafe content in T2I models by less than 20%, indicating that their generalization capabilities are insufficient.
Table 7: The performance of $\textit{Safe-Control}_{multiple}$ in maintaining benign generation for prompts from COCO 2017 val, and its comparison with baselines. Note that the downward arrow $\downarrow$ indicates that a lower score is more desirable, while the upward arrow $\uparrow$ suggests that a higher score is preferable.
@@
Table 7 displays the performance of $\textit{Safe-Control}_{multiple}$ in maintaining benign generation for prompts from MS COCO 2017 val, and its comparison with baselines. Note that the downward arrow $\downarrow$ indicates that a lower score is more desirable, while the upward arrow $\uparrow$ suggests that a higher score is preferable. We obtain two key observations from Table 4.
RQ3: Robustness Against Malicious Attacks
@@
Results. Table 8 shows the robustness evaluation of our proposed $\textit{Safe-Control}_{multiple}$ method compared to several baseline approaches—Original SD, SLD, InterpreteDiffusion, ESD, and SAFEGEN—under two malicious attacks. The evaluation metrics assess the probability of unsafe content in images generated by T2I models equipped with different defenses.
@@
Table 8: The robustness of our proposed $\textit{Safe-Control}_{multiple}$ and baselines against latest malicious attacks. We report the unsafe probability of images being generated by two attacks.